trying to help son with math homework, please help.
45 divided by 9=___ divided by 8
12 divided by ____=15 divided by 5

Found 2 solutions by Tatiana_Stebko, scott8148:
Answer by Tatiana_Stebko(1539)

1. 45 divided by 9 is 5
2. Multiply 5 by 8 =40
Therefore
45 divided by 9=40 divided by 8
.
.
.
12 divided by ____=15 divided by 5
1. 15 divided by 5 is 3
2. 12 divided by 3 is 4
Therefore
12 divided by 4=15 divided by 5

Answer by scott8148(6628)

45 ÷ 9 = 5 ___ 8 x 5 = 40 , so 45 ÷ 9 = 40 ÷ 8

15 ÷ 5 = 3 ___ 12 ÷ 3 = 4 , 12 ÷ 4 = 15 ÷ 5